# Ortho-Radial Drawing in Near-Linear Time

### v2

@@ 5 Extracting a strictly monotone cycle
Proof.
By Lemma 5.11, the facial cycle $C$ of the central face of $H$ is a strictly monotone cycle of $G$. The computation of $H$ and $C$ can be done in $O(n)$ time. ∎
We are now ready to prove Theorem 1.
See 1
Proof.
We run the $O(n\log n)$-time greedy algorithm of Lemma 4.1 for the given input $(\mathcal{R},e^{\star})$ such that $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$ for the horizontal segment $S\in\mathcal{S}_{\mathsf{h}}$ that contains $e^{\star}$. There are two possible outcomes of the algorithm. If we obtain a good sequence that covers all horizontal segments $\mathcal{S}_{\mathsf{h}}$, then we may use Lemma 3.2 to compute a drawing of $(\mathcal{R},e^{\star})$. Otherwise, the algorithm stops with a good sequence that does not cover all horizontal segments $\mathcal{S}_{\mathsf{h}}$, and no more progress can be made, in which case a strictly monotone cycle can be computed in $O(n)$ time by Lemma 5.12. ∎
@@ 6 Searching for the reference edge
In this section, we demonstrate how the drawing algorithm of Theorem 1, designed for an ortho-radial representation $\mathcal{R}$ with a fixed reference edge $e^{\star}$, can be extended to handle an ortho-radial representation $\mathcal{R}$ without a fixed reference edge. This extension incurs an additional $O(\log n)$ factor in time complexity. The core idea is to use a binary search to find a reference edge $e^{\star}$ for which $(\mathcal{R},e^{\star})$ is drawable, provided such an edge exists.
Throughout this section, we write $\ell_{C}^{e^{\star}}(e)$ to denote the edge label of $e$ with respect to $C$ using $e^{\star}$ as the reference edge. We first show the following lemma, which is essentially the same as [BNRW23, Lemma 4.2]. We still include a proof for the sake of completeness.
@@
In particular, (R3) is satisfied if and only if the intersection of $[L_{C},U_{C}]$ over all essential cycles $C$ is non-empty. In other words, (R3) is satisfied if and only if $L^{\ast}\leq U^{\ast}$, where we define $L^{\ast}=\max L_{C}$ and $\min U_{C}=U^{\ast}$, where the range of minimum and maximum is the set of all essential cycles $C$.
@@
Given that (R1)–(R3) are satisfied for the given ortho-radial representation $\mathcal{R}$, we may use a binary search to find a reference edge $e^{\star}$ such that $(\mathcal{R},e^{\star})$ is drawable, as follows. We fix any edge $e^{\prime}$ in $\overline{C_{F_{\mathsf{o}}}}$ and compute the interval $I$ defined above. We start the binary search with the interval $I$. In each step of the binary search, we let $x$ be a median of the current interval, and then we run the algorithm of Theorem 1 with a choice of the reference edge $e^{\star}$ such that $\ell_{\overline{C_{F_{\mathsf{o}}}}}^{e^{\star}}(e^{\prime})=x$.
@@
There is still one remaining issue in implementing the above approach: The algorithm of Theorem 1 requires that the reference edge $e^{\star}$ lies on a horizontal segment $S\in\mathcal{S}_{\mathsf{h}}$ with $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$. Equivalently, a choice of the reference edge $e^{\star}\in\overline{C_{F_{\mathsf{o}}}}$ satisfies this requirement if and only if $e^{\star}\in S$ such that $S$ meets one of the following requirements.
@@
We claim that for any $x\in\left[L_{\overline{C_{F_{\mathsf{o}}}}},U_{\overline{C_{F_{\mathsf{o}}}}}\right]$, we may find an edge $e^{\star}$ with $\ell_{C}^{e^{\star}}(e^{\prime})=x$ and satisfying the above requirement, so we may use this edge $e^{\star}$ as the reference edge when we run the algorithm of Theorem 1 during the binary search. Note that if $x\notin\left[L_{\overline{C_{F_{\mathsf{o}}}}},U_{\overline{C_{F_{\mathsf{o}}}}}\right]$, then the cycle $\overline{C_{F_{\mathsf{o}}}}$ is already a strictly monotone cycle under any choice of $e^{\star}$ with $\ell_{C}^{e^{\star}}(e^{\prime})=x$, so there is no need to run the algorithm of Theorem 1.
@@
Proof.
For notational simplicity, in this proof, we write $C=\overline{C_{F_{\mathsf{o}}}}$. We first start with any choice of $e^{\star}\in C$ with $\ell_{C}^{e^{\star}}(e^{\prime})=x$ and use $e^{\star}$ as the reference edge. Since $x\in\left[L_{C},U_{C}\right]$, $C$ is not a strictly monotone cycle. If $\ell_{C}(e)=0$ for all $e\in C$, then $S=C$ itself is already a horizontal segment $S\in\mathcal{S}_{\mathsf{h}}$ with $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$, so we are done. Otherwise, $C$ contains edges with labels $-1$ and $1$. We select two edges $e^{-}\in C$ and $e^{+}\in C$ in such a way that $\ell_{C}(e^{-})=-1$, $\ell_{C}(e^{+})=1$, and the length of the subpath $P$ of $C$ starting at $e^{-}$ and ending at $e^{+}$ is minimized. Our choice of $P$ implies that all intermediate edges $\tilde{e}$ in $P$ have $\ell_{C}(\tilde{e})=0$, so they form a desired horizontal segment $S$. If $e^{\star}\in S$, then we are done. Otherwise, since $\ell_{\overline{C_{F_{\mathsf{o}}}}}^{e^{\star}}(\tilde{e})=\ell_{C}(\tilde{e})=0$ for any $\tilde{e}\in S$, Lemma 6.1 implies that all edge labels remain unchanged even if we change the reference edge from $e^{\star}$ to $\tilde{e}$. Therefore, $S$ is still a horizontal segment $S\in\mathcal{S}_{\mathsf{h}}$ with $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$ even if the underlying reference edge is any $\tilde{e}\in S$. Any such an edge $\tilde{e}$ satisfies the requirement of the lemma, as $\ell_{C}^{\tilde{e}}(e^{\prime})=\ell_{C}^{e^{\star}}(e^{\prime})-\ell_{C}^{e^{\star}}(\tilde{e})=x-0=x$ by Lemma 6.1 with $e=e^{\prime}$, $e^{\star}_{1}=\tilde{e}$, and $e^{\star}_{2}=e^{\star}$. ∎
We are now ready to prove Theorem 2.
See 2
Proof.
We just need to show that for the case where the given ortho-radial representation $\mathcal{R}$ is drawable, an ortho-radial drawing realizing $\mathcal{R}$ can be computed in $O(n\log^{2}n)$ time. In this case, using Lemma 6.2, the binary search algorithm discussed above finds a reference edge $e^{\star}$ such that the $O(n\log n)$-time algorithm of Theorem 1 outputs an ortho-radial drawing realizing $(\mathcal{R},e^{\star})$. This drawing is also an ortho-radial drawing realizing $\mathcal{R}$. The number of iterations of the binary search is $O\left(\log|I|\right)=O(\log n)$, so the overall time complexity is $O(n\log^{2}n)$, as the cost per iteration is $O(n\log n)$, due to Theorem 1. ∎
@@ 7 A reduction to biconnected simple graphs
In this section, we show that, without loss of generality, we may assume that the input graph is simple and biconnected. Specifically, given a planar graph $G=(V,E)$, a combinatorial embedding $\mathcal{E}$ of $G$, and an ortho-radial representation $\mathcal{R}$ of $(G,\mathcal{E})$ satisfying (R1) and (R2), we will construct a biconnected simple planar graph $G^{\prime}=(V^{\prime},E^{\prime})$, a combinatorial embedding $\mathcal{E}^{\prime}$ of $G^{\prime}$, and an ortho-radial representation $\mathcal{R}^{\prime}$ of $(G^{\prime},\mathcal{E}^{\prime})$ satisfying (R1) and (R2) such that $\mathcal{R}$ is drawable if and only if $\mathcal{R}^{\prime}$ is drawable. See Figure 19 for an illustration of the reduction. Moreover, the reduction costs only linear time in the following sense:
@@
For the construction of the edge set $E^{\prime}$, we first add an edge between $v_{i,j}$ and $v_{i^{\prime},j^{\prime}}$ if $|i-i^{\prime}|+|j-j^{\prime}|=1$, so $X_{v}$ becomes a grid graph, see Figure 20.
Figure 20: The grid graph $X_{v}$.
The interior angles in the four 4-cycles in the grid graph, highlighted by $\star$ in Figure 20, are all set to $90^{\circ}$ in $\mathcal{R}^{\prime}$ to ensure that $X_{v}$ must be drawn as a grid. We write the four boundary paths of the grid as follows.
@@ 8 Conclusions
In this paper, we presented a near-linear time algorithm to decide whether a given ortho-radial representation is drawable, improving upon the previous quadratic-time algorithm [BNRW23]. If the representation is drawable, then our algorithm outputs an ortho-radial drawing realizing the representation. Otherwise, our algorithm outputs a strictly monotone cycle to certify the non-existence of such a drawing. Given the broad applications of the topology-shape-metric framework in orthogonal drawing, we anticipate that our new ortho-radial drawing algorithm will be relevant and useful in future research in this field.
While there has been extensive research in orthogonal drawing, much remains unknown about the computational complexity of basic optimization problems in ortho-radial drawing. In particular, the problem of finding an ortho-radial representation that minimizes the number of bends has only been addressed by a practical algorithm [NR20] that has no provable guarantees. It remains an intriguing open question to determine to what extent bend minimization is polynomial-time solvable for ortho-radial drawing. [^2] The paper [BNRW23] also lists the computation of a bend-minimized ortho-radial representation for a given plane graph as an open question. In the preliminary version (arXiv:2106.05734v1) of the same paper [BNRW23], the authors claimed this problem to be NP-hard (Theorem 7), seemingly resolving the problem. However, the result was removed in the journal version, indicating a potential issue with their reduction, so the problem remains open. To the best of our knowledge, even deciding whether a given plane graph admits an ortho-radial drawing without bends is not known to be polynomial-time solvable.
@@
There was a series of work in finding compact orthogonal drawings according to various complexity measures [BBD+22, BDBD+00, BES12, KM99, Pat01]. To what extent can the ideas developed in these works be applied to ortho-radial drawings?

### v3

@@ 5 Extracting a strictly monotone cycle
Proof 5.24.
By Lemma 5.21, the facial cycle $C$ of the central face of $H$ is a strictly monotone cycle of $G$. The computation of $H$ and $C$ can be done in $O(n)$ time.
We are now ready to prove Theorem 2.5.
See 2.5
Proof 5.25.
We run the $O(n\log n)$-time greedy algorithm of Lemma 4.1 for the given input $(\mathcal{R},e^{\star})$ such that $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$ for the horizontal segment $S\in\mathcal{S}_{\mathsf{h}}$ that contains $e^{\star}$. There are two possible outcomes of the algorithm. If we obtain a good sequence that covers all horizontal segments $\mathcal{S}_{\mathsf{h}}$, then we may use Lemma 3.3 to compute a drawing of $(\mathcal{R},e^{\star})$. Otherwise, the algorithm stops with a good sequence that does not cover all horizontal segments $\mathcal{S}_{\mathsf{h}}$, and no more progress can be made, in which case a strictly monotone cycle can be computed in $O(n)$ time by Lemma 5.23.
@@ 6 Searching for the reference edge
In this section, we demonstrate how the drawing algorithm of Theorem 2.5, designed for an ortho-radial representation $\mathcal{R}$ with a fixed reference edge $e^{\star}$, can be extended to handle an ortho-radial representation $\mathcal{R}$ without a fixed reference edge. This extension incurs an additional $O(\log n)$ factor in time complexity. The core idea is to use a binary search to find a reference edge $e^{\star}$ for which $(\mathcal{R},e^{\star})$ is drawable, provided such an edge exists.
Throughout this section, we write $\ell_{C}^{e^{\star}}(e)$ to denote the edge label of $e$ with respect to $C$ using $e^{\star}$ as the reference edge. We first show the following lemma, which is essentially the same as [barth2023topology, Lemma 4.2]. We still include a proof for the sake of completeness.
@@
In particular, 3 is satisfied if and only if the intersection of $[L_{C},U_{C}]$ over all essential cycles $C$ is non-empty. In other words, 3 is satisfied if and only if $L^{\ast}\leq U^{\ast}$, where we define $L^{\ast}=\max L_{C}$ and $\min U_{C}=U^{\ast}$, where the range of minimum and maximum is the set of all essential cycles $C$.
@@
Given that 13 are satisfied for the given ortho-radial representation $\mathcal{R}$, we may use a binary search to find a reference edge $e^{\star}$ such that $(\mathcal{R},e^{\star})$ is drawable, as follows. We fix any edge $e^{\prime}$ in $\overline{C_{F_{\mathsf{o}}}}$ and compute the interval $I$ defined above. We start the binary search with the interval $I$. In each step of the binary search, we let $x$ be a median of the current interval, and then we run the algorithm of Theorem 2.5 with a choice of the reference edge $e^{\star}$ such that $\ell_{\overline{C_{F_{\mathsf{o}}}}}^{e^{\star}}(e^{\prime})=x$.
@@
There is still one remaining issue in implementing the above approach: The algorithm of Theorem 2.5 requires that the reference edge $e^{\star}$ lies on a horizontal segment $S\in\mathcal{S}_{\mathsf{h}}$ with $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$. Equivalently, a choice of the reference edge $e^{\star}\in\overline{C_{F_{\mathsf{o}}}}$ satisfies this requirement if and only if $e^{\star}\in S$ such that $S$ meets one of the following requirements.
@@
We claim that for any $x\in\left[L_{\overline{C_{F_{\mathsf{o}}}}},U_{\overline{C_{F_{\mathsf{o}}}}}\right]$, we may find an edge $e^{\star}$ with $\ell_{C}^{e^{\star}}(e^{\prime})=x$ and satisfying the above requirement, so we may use this edge $e^{\star}$ as the reference edge when we run the algorithm of Theorem 2.5 during the binary search. Note that if $x\notin\left[L_{\overline{C_{F_{\mathsf{o}}}}},U_{\overline{C_{F_{\mathsf{o}}}}}\right]$, then the cycle $\overline{C_{F_{\mathsf{o}}}}$ is already a strictly monotone cycle under any choice of $e^{\star}$ with $\ell_{C}^{e^{\star}}(e^{\prime})=x$, so there is no need to run the algorithm of Theorem 2.5.
@@
Proof 6.4.
For notational simplicity, in this proof, we write $C=\overline{C_{F_{\mathsf{o}}}}$. We first start with any choice of $e^{\star}\in C$ with $\ell_{C}^{e^{\star}}(e^{\prime})=x$ and use $e^{\star}$ as the reference edge. Since $x\in\left[L_{C},U_{C}\right]$, $C$ is not a strictly monotone cycle. If $\ell_{C}(e)=0$ for all $e\in C$, then $S=C$ itself is already a horizontal segment $S\in\mathcal{S}_{\mathsf{h}}$ with $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$, so we are done. Otherwise, $C$ contains edges with labels $-1$ and $1$. We select two edges $e^{-}\in C$ and $e^{+}\in C$ in such a way that $\ell_{C}(e^{-})=-1$, $\ell_{C}(e^{+})=1$, and the length of the subpath $P$ of $C$ starting at $e^{-}$ and ending at $e^{+}$ is minimized. Our choice of $P$ implies that all intermediate edges $\tilde{e}$ in $P$ have $\ell_{C}(\tilde{e})=0$, so they form a desired horizontal segment $S$. If $e^{\star}\in S$, then we are done. Otherwise, since $\ell_{\overline{C_{F_{\mathsf{o}}}}}^{e^{\star}}(\tilde{e})=\ell_{C}(\tilde{e})=0$ for any $\tilde{e}\in S$, Lemma 6.1 implies that all edge labels remain unchanged even if we change the reference edge from $e^{\star}$ to $\tilde{e}$. Therefore, $S$ is still a horizontal segment $S\in\mathcal{S}_{\mathsf{h}}$ with $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$ even if the underlying reference edge is any $\tilde{e}\in S$. Any such an edge $\tilde{e}$ satisfies the requirement of the lemma, as $\ell_{C}^{\tilde{e}}(e^{\prime})=\ell_{C}^{e^{\star}}(e^{\prime})-\ell_{C}^{e^{\star}}(\tilde{e})=x-0=x$ by Lemma 6.1 with $e=e^{\prime}$, $e^{\star}_{1}=\tilde{e}$, and $e^{\star}_{2}=e^{\star}$.
We are now ready to prove Theorem 2.6.
See 2.6
Proof 6.5.
We just need to show that for the case where the given ortho-radial representation $\mathcal{R}$ is drawable, an ortho-radial drawing realizing $\mathcal{R}$ can be computed in $O(n\log^{2}n)$ time. In this case, using Lemma 6.3, the binary search algorithm discussed above finds a reference edge $e^{\star}$ such that the $O(n\log n)$-time algorithm of Theorem 2.5 outputs an ortho-radial drawing realizing $(\mathcal{R},e^{\star})$. This drawing is also an ortho-radial drawing realizing $\mathcal{R}$. The number of iterations of the binary search is $O\left(\log|I|\right)=O(\log n)$, so the overall time complexity is $O(n\log^{2}n)$, as the cost per iteration is $O(n\log n)$, due to Theorem 2.5.
@@ 7 A reduction to biconnected simple graphs
In this section, we show that, without loss of generality, we may assume that the input graph is simple and biconnected. Specifically, given a planar graph $G=(V,E)$, a combinatorial embedding $\mathcal{E}$ of $G$, and an ortho-radial representation $\mathcal{R}$ of $(G,\mathcal{E})$ satisfying 1 and 2, we will construct a biconnected simple planar graph $G^{\prime}=(V^{\prime},E^{\prime})$, a combinatorial embedding $\mathcal{E}^{\prime}$ of $G^{\prime}$, and an ortho-radial representation $\mathcal{R}^{\prime}$ of $(G^{\prime},\mathcal{E}^{\prime})$ satisfying 1 and 2 such that $\mathcal{R}$ is drawable if and only if $\mathcal{R}^{\prime}$ is drawable. See Figure 18 for an illustration of the reduction. Moreover, the reduction costs only linear time in the following sense:
@@
For the construction of the edge set $E^{\prime}$, we first add an edge between $v_{i,j}$ and $v_{i^{\prime},j^{\prime}}$ if $|i-i^{\prime}|+|j-j^{\prime}|=1$, so $X_{v}$ becomes a grid graph, see Section 7.
{sidefigure}
[ht!] [Uncaptioned image] The grid graph $X_{v}$.
The interior angles in the four 4-cycles in the grid graph, highlighted by $\star$ in Section 7, are all set to $90^{\circ}$ in $\mathcal{R}^{\prime}$ to ensure that $X_{v}$ must be drawn as a grid. We write the four boundary paths of the grid as follows.
@@ 8 Conclusions
In this paper, we presented a near-linear time algorithm to decide whether a given ortho-radial representation is drawable, improving upon the previous quadratic-time algorithm [barth2023topology]. If the representation is drawable, then our algorithm outputs an ortho-radial drawing realizing the representation. Otherwise, our algorithm outputs a strictly monotone cycle to certify the non-existence of such a drawing. Given the broad applications of the topology-shape-metric framework in orthogonal drawing, we anticipate that our new ortho-radial drawing algorithm will be relevant and useful in future research in this field.
While there has been extensive research in orthogonal drawing, much remains unknown about the computational complexity of basic optimization problems in ortho-radial drawing. In particular, the problem of finding an ortho-radial representation that minimizes the number of bends has only been addressed by a practical algorithm [niedermann2020integer] that has no provable guarantees. It remains an intriguing open question to determine to what extent bend minimization is polynomial-time solvable for ortho-radial drawing. [^1] The paper [barth2023topology] also lists the computation of a bend-minimized ortho-radial representation for a given plane graph as an open question. In the preliminary version (arXiv:2106.05734v1) of the same paper [barth2023topology], the authors claimed this problem to be NP-hard (Theorem 7), seemingly resolving the problem. However, the result was removed in the journal version, indicating a potential issue with their reduction, so the problem remains open. To the best of our knowledge, even deciding whether a given plane graph admits an ortho-radial drawing without bends is not known to be polynomial-time solvable.
@@
There was a series of work in finding compact orthogonal drawings according to various complexity measures [JGAA-595, bridgeman2000turn, JGAA-263, klau1999optimal, patrignani2001complexity]. To what extent can the ideas developed in these works be applied to ortho-radial drawings?
[…]